Use the method of forward substitutions to solve the recurrence

T(n) = 1 + 3 T(n − 1) for n ≥ 1 , T(0) = 0.
Answer #1

Answer:

T(n)
= 3T(n-1) + 1
= 3(3T(n-2) + 1) + 1
= 3(3(3T(n-3) + 1) + 1) + 1
= 3^3T(n-3) + 3^2 + 3^1 + 1
= 3^nT(1) + 3^n-1 + ... + 1
= 3^n + 3^n-1 + ... + 1
= (3^(n+1) - 1) / 2
so, T(n) = O(3^n)

Answer: T(n) = O(3^n)